free solution plz snd me



Find the particular integral solution of  d^3y/dx^3 + 4(dy/dx)=x+3cosx+ e^-2x


1
Expert's answer
2021-02-16T12:34:23-0500

Solution.

"\\frac{d^3y}{dx^3}+4\\frac{dy}{dx}=x+3\\cos\u2061{x}+e^{\u22122x}."

This is a linear differential equation of the third-order with constant coefficients. The general solution of this equation is found as the sum of the general solution "\tilde{y}" of the corresponding homogeneous equation and some particular integral solution "y_*" of the inhomogeneous equation: "y=\tilde{y}+y_*."

Сompose and solve the characteristic equation:


"\lambda^3+4\lambda=0,
\newline\lambda(\lambda^2+4)=0,
\newline
\lambda_1=0, \text{or } \lambda^2+4=0,
\newline \text{ } 
\lambda_2=2i, \text{ } \lambda_3=-2i."


Therefore, "\tilde{y}=C_1e^{\lambda_1x}+C_2e^{ax}\cos{bx}+C_3e^{ax}\sin{bx},"

where "a" - real part, and "b" - imaginary part of a complex number "\lambda_2 \text{ and } \lambda_3." So,


"\newline \tilde{y}=C_1+C_2\cos{2x}+C_3\sin{2x}."

The particular integral solution "y_*" write in the form:


"y_\u2217=y_1+y_2+y_3\u200b."

Since the right-hand side contains a "\cos{x}" , we look for a particular integral solution "y_1" ​in the form:


"y_1=A\cos{x}+B\sin{x}."

Using the method of indefinite coefficients we find that "A=0" and "B=1." So, "y_1=\sin{x}."

Since the right-hand side also contains a "e^{-2x}" , we find a particular integral solution "y_2" ​in the form:


"y_2=Ce^{-2x}."

Using the method of indefinite coefficients we find that "C=-\frac{1}{16}."

Therefore, "y_2=-\frac{1}{16}e^{-2x}."

The number is characteristic, so "y_3" is written as


"y_3=x(Dx+E)."

Using the method of indefinite coefficients we find that "D=\frac{1}{8}" and "E=0." So, "y_3=x(\frac{1}{8}x+0)=\frac{1}{8}x^2."

As follows, the particular integral solution of "\\frac{d^3y}{dx3}+4\\frac{dy}{dx}=x+3\\cos\u2061{x}+e^{\u22122x}" is


"y_*=\sin{x}-\frac{1}{16}e^{-2x}+\frac{1}{8}x^2,"

and the general solution of the same equation is

"y=C_1+C_2\cos{2x}+C_3\sin{2x}+\sin{x}-\frac{1}{16}e^{-2x}+\frac{1}{8}x^2."

Answer.

"y_*=\sin{x}-\frac{1}{16}e^{-2x}+\frac{1}{8}x^2," where "y_*" is the particular integral solution.